OpenStudy (anonymous):

Please Check My Work?!?! Having survived the meteor impact, thanks to some last minute evasive maneuvers, the Mathonauts now set their sights on their Interstellar Headquarters. The Interstellar Headquarters orbits the Earth based on the equation y^2 + x^2 = 40,000. Using the original trajectory of the ship and complete sentences, explain to the pilot how to find where the ship’s path will cross the Interstellar Headquarters’ path if the ship's original trajectory is y = x.

OpenStudy (anonymous):

lol they really stretch for these "word problems" don't they?

OpenStudy (anonymous):

So, what i did is used substitution and plugged x in for y: so i got x^2 + x^2 = 40,000 2x^2 = 40,000 so, i divided by 2 to get x^2 = 20,000 square root each side.. then i got x = 141.42 y= 141.42

OpenStudy (anonymous):

and lol yes they do :D

OpenStudy (anonymous):

i think you are being asked to solve where the line \(y=x\) crosses the circle \(x^2+y^2=40,000\)

OpenStudy (anonymous):

you have a mistake i think it is \(x^2+y^2\) not \(x^2+x^2\)

OpenStudy (anonymous):

yeah... and this is Systems of Equations, by the way...

OpenStudy (anonymous):

ooh i see, \(y=x\) you are right, i am wrong

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

sorry, you are completely correct \[2x^2=40,000\\ x^2=20,000\\ x=\pm\sqrt{20,000}\]

OpenStudy (anonymous):

don't forget the \(\pm\) there, there are two places where the line crosses the circle

OpenStudy (anonymous):

Thank you soo sooo much... :D

OpenStudy (anonymous):



OpenStudy (anonymous):

yw you did all the work

OpenStudy (anonymous):

:D you checked it for me though :D